How many photons are produced in a laser pulse of 0.340 J at 463 nm?
Answer #1

Energy f wavelength asie vielated as, E = he h+ Plancks Constant (6.626x10-34 J/s) C™ Speed of light (3x108 m/s)Ene.494 of 1 photon = 6.626 X 10-34 x 3 x 108 (463x10-9) (1 nm = 10-9m) 6.626 x 3 x 10-34+8-(-9) 463 70-04293 x 10-34+8+ 9 =